### v1

# CAIRe: Cultural Attribution of Images by Retrieval-Augmented Evaluation

Abstract
As text-to-image models become increasingly prevalent, ensuring their equitable performance across diverse cultural contexts is critical. Efforts to mitigate cross-cultural biases have been hampered by trade-offs, including a loss in performance, factual inaccuracies, or offensive outputs. Despite widespread recognition of these challenges, [^1] For example, in a well-known 2024 incident Google Gemini’s T2I diversification techniques produced offensively ahistorical images [15] an inability to reliably measure these biases has stalled progress. To address this gap, we introduce CAIRe  [^2] Available for download at github.com/siddharthyayavaram/CAIRE, a novel evaluation metric that assesses the degree of cultural relevance of an image, given a user-defined set of labels. Our framework grounds entities and concepts in the image to a knowledge base and uses factual information to give independent graded judgments for each culture label. On a manually curated dataset of culturally salient but rare items built using language models, CAIRe surpasses all baselines by 28% F1 points. Additionally, we construct two datasets for culturally universal concepts, one comprising of T2I generated outputs and another retrieved from naturally-occurring data. CAIRe achieves Pearson’s correlations of 0.56 and 0.66 with human ratings on these sets, based on a 5-point Likert scale of cultural relevance. This demonstrates its strong alignment with human judgment across diverse image sources.
1 Introduction
Figure 1: We propose CAIRe, a novel evaluation metric that assesses the degree of cultural relevance of an image, given a user-defined set of labels. Unlike existing methods that assume a definition of culture, we let the user specify the proxy of culture such as countries, religions, or cities to assess over as free-text labels.
Current text-to-image (T2I) models, despite their high-quality outputs, have serious issues in cultural representation and sensitivity. Previous research has reported that they produce culturally homogeneous outputs given under-specified prompts [24, 42] and their outputs are disproportionately biased toward Western cultures, failing to depict global diversity [44]. Further, they generate stereotypical, offensive and factually incorrect outputs when asked to be culturally inclusive [4, 8, 21, 52]. The proliferation of AI-generated content, particularly images, is reshaping our digital media ecosystem, with over 15 billion AI-generated images created since 2022.333https://wired.me/culture/ai-image/ Given this widespread adoption, it is imperative to ensure that these models are inclusive and unbiased towards users from diverse backgrounds.
Moving towards this ideal requires us to first develop robust evaluation metrics that can identify and quantify these cultural representation gaps. But how do we define culture? In the social sciences, culture is a complex concept that can refer to cultural heritage [5], social interactions [33], or ways of life [38]. It transcends basic categorizations; spanning ethnicities, communities, and even subtle neighborhood distinctions like those between upper and downtown Manhattan [6, 10], and is difficult to define concretely because it varies by context. Every individual and group lies at the intersection of multiple cultures (defined by their political, professional, religious, regional, class-based and other affiliations) and these are invoked according to the situation, typically in contrast to other groups [1]. Despite this rich and nuanced complexity, most AI research defaults to using countries as proxies for culture – indeed, this is one of the most frequently cited limitations across these same papers [62]. Cultural AI work also struggles with capturing culture’s inherently dynamic nature. While culture continuously evolves through social negotiation [35], existing benchmarks remain largely static collections of examples or facts [46, 25, 22, 28]. To address both the fluidity of cultural boundaries and their evolution over time, it is necessary to allow users to define culture on their own terms through natural language descriptions, creating a framework that can adapt as cultures themselves transform.
In this paper we introduce Cultural Attribution of Images by Retrieval (CAIRe), a family of evaluation metrics for visual cultural attribution. CAIRe evaluates images over user-defined cultural labels, by grounding them to entities and concepts in a knowledge-base (§3). Our approach takes as input an image and a list of free-text labels representing cultures of interest. It then outputs a score on a five point scale, indicating the relevance of the image to each label in the culture set (Figure 1). The framework operates in two key stages: First, it grounds the image in real-world concepts and objects, leveraging a massively multilingual and multicultural knowledge base. Second, it utilizes a vision-language model (VLM) to estimate cultural relevance in a retrieval-augmented evaluation setup, where all available information about the recognized concepts is leveraged to estimate this score. Importantly, our method is designed as a flexible framework, allowing users to integrate their preferred knowledge bases or VLMs to suit their specific needs and contexts.
All previous works which have attempted to evaluate geographical [18] or cultural diversity [24] of images assign a single country/region culture label to an image with a binary (relevant / not relevant) score. In contrast, CAIRe estimates the degree of relevance across multiple user-defined culture labels. Due to the absence of test sets with such annotations, we first construct a dataset (§5) of rare and culturally significant items, labeled using GPT-4o [36] and verified with Wikipedia. On this test set CAIRe surpasses baselines by 28% F1 points. After validating CAIRe on this dataset, we evaluate its correlation with human judgments of cultural relevance across a broad range of universal concepts. CAIRe is adept at capturing annotators’ judgments of the relevance of these images to their own cultures, with Pearson’s correlations of 0.56 and 0.66 to these human ratings over generated and natural images, respectively, demonstrating its alignment with human judgment across diverse image sources (§6). In summary, we:
  • Formalize the task of visual cultural attribution—providing graded assessments on how relevant an image is to a user-defined set of free-text culture labels.
  • Build two complementary test sets using our formalism to test visual cultural attribution methods. The first covers obscure concepts spanning a variety of culture proxies and is labeled using GPT-4o, [^4] https://huggingface.co/datasets/cmu-lti/caire-specific while the second covers universal concepts and is annotated with human judgment on a five-point scale. [^5] https://huggingface.co/datasets/cmu-lti/caire-universal
  • Introduce the CAIRe framework built using our design principles, which leverages vision-language models, massive KBs, and image-based retrieval to perform visual cultural attribution.
Figure 2: Overview of CAIRe. From an image-indexed multimodal knowledge base, we embed an input image to retrieve entities that are tied to Wikipedia articles. From the text of those Wikipedia articles and the query image, a vision-language model (VLM) generates an affinity score to each user-specified candidate culture label. A detailed description of our framework is in §3.
2 Task Formulation
We define visual cultural attribution as the task of assigning cultural relevance scores to an image, given a user-defined set of culture labels. Defining the relevance of an image with respect to a particular culture presents several challenges. An effective framework must:
  1. 1.
    Allow flexible definition of culture labels: Culture cannot always be represented by simple country or region-based labels; finer-grained community, ethnic, or social group labels must be considered. Thus, it is necessary to allow labels to be defined flexibly using natural language.
  2. 2.
    Provide graded judgments: Cultural relevance is not binary; many cultural elements are shared across different groups to varying extents.
Formally, given an input image $I$ and a set of cultural labels $\mathcal{C}=\{c_{1},c_{2},\dots,c_{n}\}$ (such as the country or religion labels shown in Figure 1), we define the scoring function $f:I\times\mathcal{C}\rightarrow[1,5]$ that outputs an integer cultural relevance score on a five-point scale. A higher score signifies higher cultural relevance of the image to the culture.
Prior work (§7) either assigns binary relevance scores to images or only allows scoring over a fixed cultural proxy label set, thereby failing to adhere to our task formulation.
3 CAIRe Metric Framework
In this section, we detail how CAIRe performs visual cultural attribution using information about similar images retrieved from a knowledge base (KB). CAIRe processes each image in two steps: (a) Visual Entity Linking (VEL), which retrieves KB entities relevant to the image using embedding-based matching [13], and (b) Cultural Relevance Scoring, which uses VLM judgments over a user-specified set of candidate culture labels. An overview of this pipeline is shown in Figure 2 and its details are below.
3.1 Visual Entity Linking (VEL)
In this step, CAIRe identifies entities and concepts present in the input image by linking them to entries in a knowledge base (KB), which provides rich textual descriptions useful for cultural relevance assessment. Our VEL design follows prior work that leverages vision-language encoders (V-L encoder) for direct image-image and image-text retrieval [47, 19].
To enable retrieval, we first index images and text from a chosen KB using a V-L encoder (see Step 0 in Figure 2). Note that the framework design allows users to integrate custom KBs and V-L encoders, making CAIRe generalizable across domains. For this work, we use the massively multilingual BabelNet graph [34] as the KB and mSigLIP 666huggingface.co/google/siglip-so400m-patch16-256-i18n [60] as the V-L encoder. BabelNet contains 6 million entities, each associated with images and text descriptions. For each entity, we compute separate mSigLIP embeddings for the linked images and texts.
These embeddings are stored in two FAISS indices [7], one for images and one for text, to enable fast similarity search. During inference, given an input image $I$, we compute its embedding $E_{I}$ and retrieve the top-$K$ most similar KB entities based on cosine similarity:
$\mathcal{N}(I)=\text{TopK}\big{(}\text{sim}(E_{I},E_{KB})\big{)}$ (1)
where $\mathcal{N}(I)$ denotes the set of top-$K$ retrieved KB entities, $E_{KB}$ are embeddings of KB images, and $\text{sim}(\cdot,\cdot)$ represents cosine similarity. We empirically choose $K=20$ in this work.
For clarity, we refer to KB entities as $e$ and candidate culture labels as $c$ throughout the remainder of this paper.
Each retrieved image corresponds to at least one KB entity, but may correspond to multiple. For example, a picture of a mango might link to both the mango and fruit nodes in BabelNet—we need to select the most precise and representative entity for cultural assessment. To disambiguate among candidates, we leverage the lemma text associated with each entity. We embed each lemma using the same V-L encoder and compute its similarity to the input image embedding. The final KB entity $e^{*}$ is selected as:
$e^{*}=\arg\max_{e\in\mathcal{N}(I)}\text{sim}(E_{I},E_{e}^{\text{text}})$ (2)
where $E_{e}^{\text{text}}$ is the text embedding of entity $e$ and $\mathcal{N}(I)$ is the set of top-$K$ KB entities retrieved in the previous step.
This two-step retrieval and disambiguation process ensures that the selected KB entity is both visually and semantically aligned with the input image. Empirical evaluation on the FOCI benchmark [14] demonstrates the effectiveness of this approach over alternative retrieval-disambiguation techniques, as discussed in Appendix §A.3.
3.2 Cultural Relevance Scoring
Once the input image is linked to relevant KB entities, we assess its cultural relevance using large language and vision-language models (LLMs/VLMs), conditioned on the candidate culture labels in $\mathcal{C}$. Specifically, we use textual descriptions from Wikipedia articles corresponding to retrieved entities, combined with the input image (when applicable), to guide the model’s judgment.
Rather than asking the model to select a single culture label, we prompt it to independently rate the relevance of the image to each candidate culture in $\mathcal{C}$ on a 1–5 scale. This formulation offers finer granularity and reflects the fact that cultural relevance is rarely binary.
Following prior work on LLMs as judges [29], which explore both numerical scoring [23] and token-likelihood methods [58], we experimented with two approaches: (i) numerical scoring and (ii) log-likelihood based scoring. In practice, numerical scoring yielded better alignment with human judgments and proved more straightforward to use. A comparison of both approaches is provided in Appendix §B.1.
In our numerical scoring setup, we prompt the model to produce an integer score between 1 and 5, based on the input Wikipedia text, the image (if available), and a rubric designed to encourage consistent expert-like assessments. An example of the prompt format is shown in Appendix §8.
To ensure reliable and well-structured outputs, we apply constrained decoding, restricting the model’s output space to valid score tokens (1, 2, 3, 4, 5). The final cultural relevance score is computed as:
$s=\arg\max_{k\in\{1,2,3,4,5\}}P(k\mid I,T,c,R)$ (3)
where $P(k\mid I,T,c,R)$ represents the model’s predicted probability of score $k$ given the input image $I$, Wikipedia text $T$, candidate culture label $c$, and rubric $R$.
This formulation produces interpretable scores that correlate well with human judgments, while remaining efficient and scalable for evaluating cultural relevance across large datasets.
(a) specific set (countries)
(b) universal-generated (countries)
(c) universal-retrieved (countries)
(d) specific set (religions)
(e) specific set (Bronze Age civilizations)
(f) specific set (cities of Indonesia)
Figure 3: Examples from the evaluation set: (a) image from the specific set depicting Azulejos. The label set of countries consists of Portugal, Spain, Brazil, Morocco, Mexico. (b) T2I-generated image using the prompt “A realistic photo of a ritual in India,” representing the universal-generated subset. (c) image retrieved from DataComp 1B, using the text query “A realistic photo of greetings in Nigeria”, illustrating the universal-retrieved subset. (d-f) additional examples from the specific set corresponding to cultural proxies religion, Bronze Age civilizations, and cities of Indonesia, (Buddhism, Sumer, and Magelang respectively).
4 Test Set Curation
Given our formulation of visual cultural attribution as requiring both dynamic definitions of culture and graded relevance judgments (§2), it is crucial to evaluate models on a test set that reflects these criteria. However, existing datasets for assessing cultural and geographic diversity focus exclusively on single, binary country-level annotations [24, 18]7), which are insufficient for our task. To address this gap, we construct two complementary test sets—specific and universal—designed to reflect CAIRe’s desiderata.
The specific set consists of real-world images depicting culturally specific, and often obscure, entities. This set enables us to assess whether CAIRe can reliably capture rare and fine-grained cultural elements. To create labels for this set, we use GPT-4o to assign relevance scores across a diverse set of cultural proxies that extend beyond simple geographic categories, directly supporting desideratum 1 from §2.
The universal set focuses on culturally universal concepts (e.g., food, ceremony) and is annotated by 200 annotators from 10 different countries. It includes two splits: generated, containing text-to-image (T2I) generated images, and retrieved, containing natural images sourced to represent each concept. Annotators rated each image on a five-point scale based on its cultural relevance to their own country. Because annotator metadata includes only country-level information, we aggregate these graded judgments across countries. This design satisfies desideratum 2 from §2, by capturing varying degrees of relevance across diverse cultural perspectives.
4.1 specific Concept Test Set
The specific set contains 68 concepts [20] produced using GPT-4o. We prompt GPT-4o to provide a list of concepts that are rare, yet significant to different proxies of culture, such as regions, religions, festival, ethnicities, philosophies, etc., as laid out in Adilazuarda et al. [1]. This produces a diverse set of entities across many candidate culture label sets, including countries of the world, ethnicities of Africa, states of India, cities of Indonesia, world religions, Native American tribes, and Bronze Age civilizations. Given one proxy, like cities of Indonesia, an image can have multiple gold labels indicating that its relevant to multiple cities of Indonesia. The complete set is provided in Appendix §C.6. Each generated concept and its labels are validated through an affirmative match to an existing Wikipedia article.
Next, we manually collect CC-licensed images from the web for each concept, ensuring that there is no exact match with images in the KB, [^7] The dataset will be open sourced along with the code to avoid test set contamination. Samples from this test set with its respective culture proxies is shown in Figure 3.
Metric:
Since GPT-4o does not reliably produce 1–5 numerical ratings, and assigning graded relevance scores is infeasible for many of these rare and culturally specific entities—especially when some cultural groups may be unfamiliar or poorly represented—we instead formulate this as a multi-label classification task. For each image, the goal is to predict the set of culture labels to which the concept is relevant. Performance is measured using F1 score and Precision, computed across all culture labels. This formulation allows us to directly assess whether CAIRe can accurately identify culturally relevant labels for obscure or fine-grained entities. We report both F1 scores and Precision for CAIRe and baseline models in §5.
4.2 universal Concept Test Set
The curated specific test set above includes multi-cultural annotations for each concept, but doesn’t allow us to measure the accuracy of CAIRe in estimating the degree of culture relevance, and how closely it matches human judgment. Hence, we curate a list of universal concepts that are common across multiple cultures (like breakfast, weddings, rituals, etc.) and collect both natural and generated images for selected concepts. Using these images, we run a human rating study on Prolific, [^8] https://www.prolific.com/ where we collect cultural relevance ratings for each image on a scale of 1 to 5. Our data curation process is detailed below:
  • Selecting Concepts: We adopt 20 visually depictable and culturally universal concepts [^9] From a set of 298 human universals available here: https://condor.depaul.edu/~mfiddler/hyphen/humunivers.htm provided by Bhatia et al. [3] (full list in Appendix §C.2). This is to ensure that annotators across diverse geographies are equally competent in understanding these concepts and judging their cultural relevance.
  • Selecting Regions: We recruit annotators from China, India, United States, Brazil, Nigeria, Russia, Mexico, Egypt, Germany, and Indonesia, geographically diverse countries that have many available annotators.
  • Generated data (universal-generated): We use Stable Diffusion 3 Medium (SD3) [11] to generate images. We use a simple prompt format: A realistic photo of CONCEPT in COUNTRY, for all combinations of concepts and countries listed above, producing 200 prompts (20 concepts × 10 countries) and 200 images per model.
  • Natural data (universal-natural): We obtain mSigLIP embeddings for Datacomp-1B [12] and retrieve the top-8 images for the same 200 prompts used in universal-generated. We manually select a single high-quality image from the top-8 images, since retrieval may be noisy for certain prompts.
Human Annotation:
We evaluate 400 images (200 generated from SD3 (universal-generated) and 200 retrieved from DataComp-1B (universal-natural)). Each image receives a cultural relevance rating on a 1-5 Likert scale by 20 annotators each, across the 10 countries mentioned above. The definition of each rating is shown in Appendix C.3 (Table 12). Each image is assigned a cultural relevance score for each country by averaging the Likert scores given by annotators in said country.
Metric:
We compute Pearson’s correlation coefficient between the absolute cultural relevance scores assigned by human raters and those predicted by our tool for each image. This metric quantifies the degree of linear association between the predicted scores and human ratings, where a high correlation indicates that our tool not only ranks countries accurately but also assigns scores that align closely with human perceptions of cultural relevance.
5 Experimental Setup
Since CAIRe is the first work proposing a metric for assessing the degree of cultural relevance of an image, we benchmark its performance against general-purpose V-L encoders and VLMs. We set up these baselines as follows:
5.1 Baseline 1: Vision-Language Encoders
The first baseline uses embeddings of the input image and a set of textual probe prompts using either CLIP ViT-B/32 [40] or mSigLIP [60] to assign a culture label.
The probe prompts are constructed from the user-defined set of candidate culture labels and a set of five relevance level phrases by combining them into the string This image is {relevance level} to {culture candidate}. Here, relevance level is one of: “Not relevant”, “Minimally relevant”, “Somewhat relevant”, “Relevant”, “Highly relevant”, following the template shown in Table 12.
We compute the cosine similarity between the input image and each of the probe prompt embeddings, selecting the label with the highest similarity score as the final rating.
5.2 Baseline 2: Vision-Language Models
We prompt three recent and popular vision-language models (VLMs) to provide a graded score on cultural relevance: Llama-3.2-11B-Vision-Instruct [16], Qwen2.5-VL-7B-Instruct [54], and Pangea-7B-hf [59]. The prompt includes the target culture along with a description of the relevance levels, as defined in Table 12. We encourage the model to follow chain-of-thought reasoning to obtain best results. The full prompt template is provided in Figure 8.
5.3 Our Metric: CAIRe
We test LLaMa-3.2-11B, Qwen2.5-VL-7B, and Pangea-7B-hf [16, 54, 59] as different LLM/VLM alternatives in the cultural relevance scoring step of CAIRe3.2).
In the experiments with LLMs, we simply provide the text retrieved in the entity linking (VEL) stage (§3.1) as context to the model, but for VLMs, we can provide both the input image and the retrieved entity information. We experiment with two possible ways of augmenting the LLMs and VLMs with additional retrieved context:
  • Wikipedia Text Augmentation: We incorporate the full Wikipedia page text corresponding to the best matching entity identified through lemma disambiguation.
  • Top-K Entity Title Augmentation: Instead of detailed descriptions, we provide only the Wikipedia titles of the top 20 retrieved entities. This allows us to test whether models can leverage their own internal knowledge when provided with minimal cues in the form of concept names. This strategy also reflects the fact that retrieval accuracy at the top-20 level is often higher than at the top-1, and thus may offer useful supplementary information.
5.4 Evaluation
CAIRe and all our baselines, produces a cultural relevance score on a five-point scale, as desired in §2. However, for the specific set, we simply have multi-label binary relevance labels, as explained in §C.6. Thus, to evaluate performance on the specific set we convert all metrics’ outputs into binary judgments by assigning scores greater than 3 on the five-point scale (relevant, highly relevant) to positive label 1, and the others 0. We report Precision & F1 scores for each metric based on these in Table 1. We also conduct binary classification at thresholds from 2–5 and observe that CAIRe consistently outperforms all baselines across these thresholds, as detailed in C.7.
For the universal human-labeled test set where we collect ratings on a scale of 1-5, we directly measure Pearson’s correlations between the baseline’s score and average human ratings, which can be found in Table 2.
6 Results and Analysis
Model Img. Wiki. Top-20 Prec. F1 $\Delta$${}_{\textsc{CAIRe}(F1)}$
Vision-Language Encoders (Baseline-1)
CLIP 5.9 3.3 -
mSigLIP 7.7 12.8 -
Vision-Language Models (Baseline-2)
Llama-3.2-11B-Vis.-Ins. 38.6 40.8 -
Qwen2.5-VL-7B-Ins. 37.4 41.0 -
Pangea-7B-hf 12.7 20.3 -
CAIRe
Llama-3.2-11B-Vis.-Ins. 51.1 47.4 (+6.6)
46.5 47.9 (+7.1)
[…]
58.5 52.1 (+11.1)
[…]
64.5 54.8 (+13.8)
[…]
20.6 29.0 (+8.7)
Table 1: Precision and F1-scores on the specific set. $\Delta_{\textsc{CAIRe}(F1)}$ represents the improvement in F1-score that CAIRe provides over the naive baseline with that same LM. Img. indicates the input image, Wiki. represents the Wikipedia content of the top-matched entity, while Top-20 refers to the names of the top-20 matched entities.
Model Brazil China Egypt Germany India Indonesia Mexico Nigeria Russia USA Avg.
(N) (G) (N) (G) (N) (G) (N) (G) (N) (G) (N) (G) (N) (G) (N) (G) (N) (G) (N) (G) (N) (G)
Vision-Language Encoders (Baseline-1)
CLIP -0.04 0.38 0.04 0.16 0.09 0.07 -0.17 -0.09 -0.05 0.12 0.28 0.33 -0.06 0.07 -0.02 0.20 -0.10 0.08 -0.02 0.11 -0.02 0.11
mSigLIP 0.07 -0.02 -0.01 -0.15 0.22 0.08 0.15 -0.04 -0.13 -0.01 0.16 -0.20 0.09 0.05 0.20 -0.02 0.15 0.21 0.12 0.03 0.10 -0.01
Vision-Language Models (Baseline-2)
LLaMA3.2-V 0.54 0.18 0.57 0.37 0.31 0.29 0.65 0.27 0.59 0.34 0.56 0.33 0.45 0.11 0.68 0.56 0.57 0.01 0.52 0.28 0.55 0.31
Qwen2.5-VL 0.59 0.51 0.59 0.37 0.54 0.52 0.56 0.54 0.76 0.72 0.64 0.60 0.52 0.39 0.62 0.61 0.69 0.33 0.66 0.63 0.61 0.52
Pangea-7B 0.56 0.39 0.60 0.50 0.54 0.58 0.62 0.52 0.72 0.66 0.56 0.47 0.50 0.40 0.46 0.60 0.61 0.39 0.64 0.59 0.58 0.51
CAIRe
Llama-3.2-V 0.64 0.47 0.61 0.30 0.52 0.49 0.70 0.58 0.69 0.70 0.65 0.44 0.57 0.33 0.47 0.49 0.74 0.53 0.65 0.61 0.63 0.49
Qwen2.5 0.57 0.41 0.66 0.34 0.22 0.43 0.48 0.34 0.62 0.67 0.67 0.47 0.55 0.36 0.42 0.44 0.66 0.29 0.48 0.42 0.53 0.42
Qwen2.5-VL 0.67 0.63 0.68 0.53 0.60 0.50 0.65 0.58 0.74 0.68 0.70 0.53 0.58 0.46 0.61 0.55 0.74 0.56 0.59 0.62 0.66 0.56
Pangea-7B 0.70 0.49 0.64 0.45 0.61 0.56 0.62 0.51 0.52 0.65 0.58 0.55 0.59 0.39 0.64 0.64 0.61 0.47 0.61 0.47 0.61 0.52
Table 2: Country-wise Pearson’s correlation with human judgment. (N) represents natural data, while (G) represents generated data. The highest and lowest country-wise values are bold and underlined, respectively. A detailed analysis of the results is provided in §6. Models used are LLaMA-3.2-11B-Vision-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-VL-7B-Instruct, and Pangea-7B.
Table 1 presents the Precision & F1 scores for the CAIRe metrics and baselines on the specific test set. Table 2 presents the Pearson correlation coefficients of each model against the human-labeled cultural relevance judgments in the universal set. Our analysis of CAIRe’s performance against the baselines answers six questions:
Does CAIRe correctly attribute specific entities to their respective cultures?
The specific set checks this using particularly difficult real-world example concepts. Table 1 shows that CAIRe consistently outperforms the open-source LM and VLM baselines at this binary cultural relevance classification task. Among all VLMs (Baselines-2), Qwen2.5-VL-7B-Instruct achieves the highest F1-score of 41.0 without additional context, representing the strongest baseline result. Integrating this VLM into CAIRe significantly improves performance (by 28 F1 points!).
How strongly does CAIRe correlate with human opinions?
The universal set images provide natural subjective annotator opinions of an image’s relevance to their own culture. Thus, correlation between metric judgments and human labels over this set characterizes how well CAIRe models these opinions. These results are presented in Table 2. The strongest CAIRe metric on the specific set, Qwen2.5–VL-7b-Instruct, is also most performant on the universal sets, achieving Pearson’s correlations of 0.56 and 0.66 on average across all countries. Note that the performance difference between the VLM baselines and CAIRe are not as pronounced on this set—this may be due to the relative commonality of concepts present in the universal images to the specific ones (Figure 3).
Does CAIRe’s model of human opinions generalize across diverse cultures?
Most columns in Table 2 contain the correlation scores averaged over all concepts within specific countries. The lowest and highest values for each model are marked using underlines and bold respectively.
Generally, CAIRe is more performant for India, Germany and Russia compared to Mexico, Egypt and Nigeria. This can be attributed to distributional inequalities in mSigLIP’s training data and how well-populated the KB is for all these regions. This may also be a result of disparities in the representativeness of the presented images in the test set. We also see that competent VLMs like Qwen display lesser variation across countries (0.5-0.7), while LLaMa displays higher variance. Overall, this suggests that with better models and larger KBs, CAIRe can consistently improve with time.
How effective is CAIRe in visual entity linking, i.e., matching images to the right entities?
The VEL system successfully matches 86% of instances in the specific set, to their exact labeled concepts, as verified through manual evaluation. In instances where an exact match is not achieved, the retrieved concepts often exhibit semantic or cultural similarity to the query image, with only a minimal number of cases being entirely incorrect. We present a few of these examples in §C.1, with most inaccuracies arising from concepts that are absent in the knowledge base.
To estimate an upper bound on how much VEL performance impacts the overall F1 score in the case of the specific set, we incorporate gold-standard Wikipedia pages as context in our cultural relevance scoring, allowing us to quantify retrieval-induced loss (detailed results in §C.5: Table 10).
Which variant of CAIRe performs best?
On the specific set, Qwen2.5-7B-Instruct achieves the highest performance, narrowly outperforming the VL variant despite not using the image. One possible explanation is that when dealing with difficult entities, Qwen-VL may sometimes conflict with retrieved content, leading to errors. In contrast, Qwen2.5-7B-Instruct, which relies solely on retrieved text, avoids this issue and thus achieves slightly better results.
Additionally, models augmented with the full Wikipedia page of the top entity generally outperform those provided only with the titles of the top 20 retrieved entities. While supplying multiple entity titles can help the model form a broader view, it may also introduce noise and conflicting signals, which can hurt performance on fine-grained cultural judgments.
For the universal set, Qwen2.5-VL-7B-Instruct consistently achieves the highest weighted average Pearson correlation across cultures for both natural and generated images, while also exhibiting the lowest variance across countries—indicating clear superiority among the tested VLMs.
Overall, the Qwen family of models delivers the strongest results with CAIRe, showing a 24–27% absolute F1 improvement over native prompting when augmented with retrieved knowledge.
How does performance vary between natural and generated images?
Per Table 2, it’s clear that CAIRe is more effective in evaluating the cultural relevance of natural images as compared to generated ones. This is expected since arises our KB comprises exclusively natural images and mSigLIP is also trained on naturally-occurring data. Furthermore, the system demonstrates strong performance on the challenging specific set comprising of natural images, reinforcing this observation. In the case of generated images, the depicted entities do not always correspond to real-world concepts or objects, even when the model is explicitly prompted to generate “realistic” images. However, as generative model capabilities advance, both in terms of realism and in accurately depicting real-world entities, this performance gap is expected to diminish.
7 Related Work
Culturally-diverse image datasets
Several studies have explored the need for culturally diverse image datasets to better assess biases in text-to-image (T2I) models. Recent work [57] demonstrated that incorporating culturally and linguistically diverse data in training can enhance the fairness and accuracy of visual representations. The GeoDE dataset [41] contains 61,940 images of common objects from geographically diverse regions to evaluate object recognition systems otherwise trained on western-centric web-scraped images. Jha et al. [21] developed ViSAGe, a dataset of T2I generated outputs, human-annotated for visual stereotypes. Khanuja et al. [26] introduce a new task and test set to evaluate whether image-editing models are capable of localizing images for a target culture. Finally, Bhatia et al. [3] create a benchmark that introduces two challenging tasks to test for cultural inclusion in vision-text models: retrieval across universals and cultural visual grounding. Unlike most of these works that have treated geographical regions as a proxy for culture and assigned single culture labels to each image, our datasets include multi-culture labels with a human rating on how relevant an image might be to each culture in the set.
Evaluating for fairness and diversity
Prior work has attempted to assess biases and fairness in T2I model outputs. DIG In [18] and Decomposed-DIG [48] evaluate whether T2I models generate geographically diverse outputs for single-object prompts, revealing representational disparities. Basu et al. [2] highlighted the Western-centric bias in these models, while Ventura et al. [49] examined embedded cultural perspectives, advocating for improved evaluation frameworks.
CUBE [24] assesses cultural diversity using a manually curated reference set limited to eight countries and three domains with 300k artifacts. In contrast, our framework leverages a knowledge base spanning 6M concepts, enabling broader assessments. Unlike prior work that primarily evaluates batch-level diversity and focuses only on generated images, our approach assesses cultural relevance at the individual image level and works with both generated and natural images. Finally, while most existing evaluations assign a single country label per image, we introduce a more flexible and inclusive method, allowing multiple cultural proxies to define and assess cultural relevance.
Work on multilingualism in text-to-image models overlaps with themes of cultural evaluation. Saxon and Wang [44] introduced an evaluation of cross-lingual capabilities of a slate of T2I models using image similarity over cultural universals. This line of work has faced challenges for a lack of techniques like CAIRe to perform cultural attribution [43, 45].
Binary v/s Graded Judgements in Cultural Contexts
: Prior work across education, communication, and psychology consistently shows that cultural relevance is best understood as a matter of degree rather than a binary condition. Binary judgments oversimplify the subjective and context-dependent nature of cultural fit, potentially masking ambivalence or partial connections [51]. Empirical studies demonstrate that graded scales more effectively capture this nuance. For instance, Ebe [9] showed that elementary students rated storybooks for cultural relevance using a rubric, and those receiving higher relevance ratings led to better comprehension and engagement. Similarly, in health communication, Vereen et al. [50] found that Black women rated culturally tailored visual advertisements significantly higher on cultural relevance than generic versions, using a 7-point Likert scale, though both ads conveyed identical health information. Narrative media research also supports this approach: [32] reported that Latino viewers rated culturally resonant videos as more culturally similar to themselves, which corresponded to stronger identification with the characters. Finally, classic advertising research by Whittler [55] found that participants rated advertisements featuring spokespeople of their own race as more relatable, with these higher identification ratings predicting more favorable attitudes toward the ad. Together, these studies affirm that cultural relevance judgments are inherently graded, and that ordinal scales better capture varying levels of perceived fit and resonance. For tasks like estimating the cultural relevance of images, which demand modeling subtle human perceptions across diverse cultural identities, graded scoring (such as a 1–5 relevance scale) is therefore far more appropriate and informative than binary labels.
8 Conclusion
From our formulation of the visual cultural attribution task, we have produced two test sets that capture important desiderata: specific accuracy over rare entities and faithful modeling of human opinions over diverse cultures.
Our CAIRe family of visual cultural attribution metrics is the first framework in this space that permits assessments over free-text culture labels (open vocabulary) and provides numerical scores rather than binary judgments. By leveraging a massively multilingual, multimodal knowledge base and state-of-the-art vision-language encoders, CAIRe can account for rare entities, and using those retrieved KB entries as input to LMs for final judgment, we robustly provide these numerical scores. Our results demonstrate that CAIRe beats naive LM-based approaches on both test sets, and satisfies our need for a robust visual cultural attribution tool.
With CAIRe, we significantly expand the scope of concepts we can evaluate for cultural diversity, since we use the KB images as our reference set for grounding. Given its modularity and leverage of state-of-the-art LMs, CAIRe can continue to evolve going forward and enable research into cultural issues in multimodality.
9 Limitations & Ethical Considerations
Biases from using Wikipedia Content.
Wikipedia is a valuable resource for training language models due to its breadth and structured content, but it also introduces biases. Its coverage reflects the interests of its editor community, leading to variations in topic emphasis and linguistic framing. Additionally, its reliance on verifiable sources can favor well-documented perspectives over emerging viewpoints. These factors influence LM outputs, reinforcing the need for careful interpretation.
Potential Reinforcement of Cultural Stereotypes.
While CAIRe aims to provide a nuanced evaluation of cultural relevance, it relies on knowledge bases and vision-language models that may contain biases. If not carefully calibrated, the framework could inadvertently reinforce existing stereotypes by overemphasizing certain cultural elements while under-representing others.
Subjectivity in Cultural Attribution.
Defining cultural relevance is inherently subjective, and different users may have differing perspectives on whether an image aligns with a particular culture. While CAIRe allows for flexible cultural definitions, the subjectivity in user-defined labels and scoring necessitates caution in interpreting results, especially in sensitive contexts.
10 Contributions
We list author contributions loosely following the CRediT author statement. [^10] https://www.elsevier.com/researcher/author/policies-and-guidelines/credit-author-statement
AY: Methodology (co-lead), Software (co-lead), Resource Development, Writing.
SY: Methodology (co-lead), Software (co-lead), Resource Development, Writing.
SK: Conceptualization (co-lead), Methodology (co-lead), Software, Validation and Analysis, Resources, Data Curation (co-lead), Writing (co-lead), Supervision (co-lead)
MS: Conceptualization (co-lead), Data Curation (co-lead), Acronymography, Methodology, Writing, Resources
GN: Supervision (co-lead), Funding Acquisition (lead), Writing.
11 Acknowledgments
We are especially grateful to Shaily Bhatt for her valuable contributions during the early stages of brainstorming for this project. We also thank the following individuals for their insightful feedback on drafts of this paper, as well as their comments during talks and presentations: members of NeuLab, Akhila Yerukola, Fernando Diaz, Danish Pruthi, Abhipsa Basu, Nithish Kannen, Partha Talukdar, and Manish Gupta.
This work was supported in part by grants from Google, the Defense Science and Technology Agency Singapore, the National Science Foundation Graduate Research Fellowship Grant No. 1650114, and the UCSB Center for Responsible Machine Learning.
Appendix A Details on Visual Entity Linking
A.1 Off-the-shelf Object Detection
In our approach to Visual Entity Linking (VEL), we initially employed a straightforward methodology by leveraging off-the-shelf object detection models to identify entities within images. We utilized Detectron 2 [56] from Facebook research along with YOLOv10 [53]. The underlying premise was to recognize object names through these detection tools and subsequently map them to corresponding knowledge base (KB) identifiers, akin to traditional textual entity linking. To this end, we experimented with state-of-the-art object detection models capable of recognizing fine-grained object categories in both open and closed-world settings, where KB identifiers served as predefined classification labels. COCO [30] and LVIS [17] classes are 2 of the most commonly used sets of classes. Additionally, we utilized Vision language models to detect entities. Specifically, we used LlaVA-NeXT [31] and Idefics2-8B [27].
However, these methods yielded suboptimal performance, particularly for long-tail entities, a challenge commonly observed in object detection [39]. Given that culturally rare or niche objects often fall within the long-tail distribution, this approach was deemed impractical for robust entity linking. [^11] A comparative analysis of the outputs produced by various object detection models and VLMs utilized in our study can be found here: https://hub.zenoml.com/project/954d212c-eca6-4ad6-a461-7cb623641732/Object Captioning. We present the results for 5 different systems; Using Detectron2, yolo10x, LlaVA-NeXT, Idefics2-8B, and an Entity-Attribute-Relation (EAR) pipeline. The results are presented on the “transcreation” dataset introduced by [26].
A.2 Encoder Models for Building KB Index
(a) Query Image
(b) mSigLIP (0.923)
(c) CLIP (0.918)
(d) DINOv2 (0.850)
Figure 4: Retrieval Comparison Across Encoders
We evaluated several image-text encoder models to compute both image-image and image-text similarity. After qualitatively analyzing the outputs from different variants, we found that mSigLIP performed the best. Its unique loss function improves image-text alignment, making it superior to other encoders such as CLIP [40] and DINOv2 [37] for our task. We conduct a qualitative analysis of retrieval performance using a small subset of images. This subset comprises all fruit images from BabelNet, with each encoder provided the same set of potential matches for retrieval. Figure 4 presents a comparative evaluation of retrieval performance across different encoders, demonstrating mSigLIP’s superior semantic image understanding and similarity scores.
A.3 Visual Entity Linking
Datasets IN-Food IN-Plant IN-Animal IN-Artifact
Methods @1 @5 @10 @20 @1 @5 @10 @20 @1 @5 @10 @20 @1 @5 @10 @20
Idefics-2B 11.58 10.95 10.72 11.91 4.60 4.20 4.68 5.03 8.51 6.86 8.30 9.10 7.48 8.19 8.82 9.43
Lemma (V-T) 23.65 50.65 57.93 62.07 16.15 42.14 51.30 57.34 22.47 52.45 62.64 70.11 11.61 22.02 25.51 27.88
Freq-Based (V-T) 22.71 45.79 52.02 57.85 26.32 46.21 53.24 56.24 31.84 56.55 63.43 68.52 10.34 19.99 23.16 25.98
Lemma (T) 4.68 13.69 18.02 22.78 3.20 9.13 12.82 17.10 5.67 15.66 21.69 28.17 1.63 4.15 5.65 7.22
Gloss (T) 3.32 8.13 11.54 15.16 2.39 7.32 10.82 15.17 4.15 12.46 17.66 23.73 1.09 2.85 3.80 4.97
Wikipedia (T) 8.40 20.48 25.93 31.83 4.99 14.52 20.50 27.83 6.84 20.76 28.60 37.20 4.38 10.72 13.84 17.03
Table 3: Retrieval accuracy at different thresholds across the ImageNet datasets for various disambiguation methods. The highest accuracy for each dataset under each setting is highlighted.
In this section, we detail further approaches to perform visual entity linking.
Image Query + Text Keys: We compute the similarity between input image embeddings and KB text embeddings to retrieve the top-20 entities, ranking them by similarity score. We build three different text-embedding indices to individually search over for top-matching keys, given the query image. The first comprises of lemma embeddings of each entity, the second comprises of gloss embeddings of each entity, [^12] https://www.babelnet.org/synset?id=bn:00015267n&orig=dog&lang=EN and the third comprises of Wikipedia page embeddings. [^13] https://www.tensorflow.org/datasets/community_catalog/huggingface/wikipedia
Image Query + Image Keys: We experiment with bypassing text-based disambiguation and rank each KB ID based on its occurrence frequency among the top-20 retrieved images, a method we refer to as frequency-based matching.
All of our methods are summarized below:
  1. 1.
    Lemma (T): Image-to-text similarity with KB lemmas.
  2. 2.
    Gloss (T): Image-to-text similarity with KB glosses.
  3. 3.
    Wikipedia (T): Image-to-text similarity with Wikipedia content.
  4. 4.
    Lemma (V-T): Vision-based similarity followed by text matching.
  5. 5.
    Frequency-Based (V-T): Most frequently occurring IDs among retrieved images.
To identify the best-performing method among the above alternatives, we conduct a quantitative evaluation using the Fine-grained Object Classification (FOCI) benchmark [14]. The FOCI benchmark includes nine datasets, with four subsets of ImageNet (IN-Food, IN-Plant, IN-Animal, IN-Artifact). We select this subset from the FOCI benchmark as we are able to map each ImageNet ID to the corresponding BabelNet ID using the structure of ImageNet-21k. This is enabled by ImageNet being built using WordNet categories and hierarchy. Dataset statistics are listed in Table 4. Each category has 10 associated images.
Dataset No. of Categories
IN-Food 563
IN-Plant 957
IN-Animal 1314
IN-Artifact 2630
Table 4: ImageNet dataset Statistics
FOCI is a rigorous benchmark that reframes image classification as a series of multiple-choice questions (MCQs). Leveraging CLIP-based similarity, it constructs challenging answer choices by selecting visually and semantically similar instances that are frequently confused with one another. We conduct our evaluation in an open-ended setting, where the model is provided only the input image, without predefined answer choices, making the task significantly more challenging.
Our evaluation aims to:
  • Benchmark our system against state-of-the-art VLMs for VEL.
  • Compare the effectiveness of various retrieval and disambiguation techniques.
  • Identify retrieval performance thresholds beyond which VEL performance saturates (e.g., Recall@X).
To enable quantitative evaluation, we map each category in the IN datasets to BabelNet entities and consider a prediction correct if it retrieves the exact BabelNet ID. This facilitates a direct comparison of different disambiguation strategies. Additionally, by evaluating retrieval at varying levels, we can determine the point at which performance saturates and establish the optimal number of retrieved entities per example.
Results on the FOCI benchmark shown in Table 3 in an open setting reveal that image-to-image retrieval combined with image-text disambiguation (Lemma (V-T); Gloss (V-T); Frequency-based (V-T)) consistently surpasses direct image-text retrieval methods (Lemma (T); Gloss (T); Wikipedia (T)). We outperform all state-of-the-art VLMs presented as baselines in the FOCI work itself, demonstrating the robustness of our framework in generally identifying objects belonging to the long tail and linking them to appropriate KB entities.
A.4 VEL Example
Figure 5: VEL Query Image:
Pysanka (Slavic Decorated Egg)
[…]
Figure 6: Top-5 Image Retrieval Results for Query Image.
This section aims to provide a detailed example of the best performing entity linking techniques, illustrating various disambiguation techniques along with their respective variations.
We begin by retrieving the top 20 images from our knowledge base (KB) based on the query image shown in Figure 5. Figure 6 illustrates the 5 highest similarity matches.
Note: While we present step-by-step results using only 5 retrieved images for clarity, our actual framework utilizes 20 images.
BabelNet ID Lemma
bn:00068196n Romania
bn:00029497n Easter
bn:00078872n Ukraine
bn:00538675n Folklore of Romania
bn:02889635n Etymology of Ukraine
bn:03096581n Pysanka
bn:00029503n Easter egg
Table 5: Unique BabelNet IDs and their lemmas.
Each retrieved image in Figure 6 corresponds to a list of BabelNet IDs as shown in Table 5. We extract and aggregate these IDs, retaining only the unique ones before proceeding to the disambiguation stage. The unique IDs and their corresponding lemmas are presented below:
After aggregation, these IDs are treated uniformly, irrespective of their initial image similarity. Our approach incorporates three primary disambiguation techniques.
lemma matching: We compute the cosine similarity between the query image embeddings and the lemma text embeddings for each BabelNet ID in the retrieved set. The IDs are ranked based on their similarity scores. Table 6 presents the top-5 ranked IDs alongside their corresponding lemmas and similarity scores. As observed, the highest-ranked lemma aligns with the gold label of the query image.
BabelNet ID Lemma Score
bn:03096581n Pysanka 0.5243
bn:00029497n Easter 0.5161
bn:00029503n Easter egg 0.5157
bn:00538675n Folklore of Romania 0.5108
bn:00068196n Romania 0.5037
bn:00078872n Ukraine 0.4985
bn:02889635n Etymology of Ukraine 0.4942
Table 6: Ranked BabelNet IDs
BabelNet ID Lemma Frequency
bn:03096581n Pysanka 13
bn:00029503n Easter egg 10
bn:00029497n Easter 6
bn:00078872n Ukraine 2
bn:00068196n Romania 1
bn:00538675n Folklore of Romania 1
bn:02889635n Etymology of Ukraine 1
Table 7: Frequency ranked BabelNet IDs
gloss matching: In this method, we utilize the gloss associated with each BabelNet ID. The gloss is a short definition or description of the entity or concept. This method follows the same approach as lemma matching, where we rank IDs according to gloss-image similarity. In this particular case, this method yields a ranking identical to ’lemma matching’.
frequency matching: This method ranks BabelNet IDs based on their frequency of occurrence within the ID lists associated with the retrieved images. The results of this approach are presented in Table 7, which reports the frequencies computed over a set of 20 retrieved images.
Appendix B Further CAIRe design considerations
B.1 Alternative Cultural relevance scoring techniques
Log-likelihoods:
In this approach, we opt to directly use the LM’s token log-likelihoods corresponding to each candidate culture label. Given an input prompt consisting of the retrieved textual information about an entity in the image, we prompt the LM with a completion in the form of: “This text is relevant to [culture label]”. We estimate the relevance of each culture label directly using the LM head log-likelihood of the completion. In the case of VLMs, we also provide the input image as additional conditioning context in calculating the completions’ log likelihoods.
We use the notation $\mathcal{L}_{m}(y,x)$ to denote the negative log likelihood of tokens $y$ from model $m$ conditioned on context $X$, or
$\mathcal{L}_{m}(y,X)=-\log P_{m}(y\mid X)$ (4)
To compute the log-likelihood of a culture label $c$, we construct context $X$ containing retrieved documents $D$, the aforementioned “This text is relevant to” prompt $p$, and input image $I$ to compute
$\textsc{CAIRe}(I,c)=\mathcal{L}_{m}(y,(D_{i},I,p))$ (5)
One problem with this approach is that models $m$ predict different base rates $\mathcal{L}_{m}(c_{i},(\emptyset,\emptyset,p))$ for different culture symbols when conditioned on the prompt alone. This introduces a bias where, for example, models will systematically prefer “relevant to Australia” over “relevant to Suriname” hindering consistent comparison of attribution scores between cultures.
To mitigate this, we apply an affine debiasing method using base rates and hyperparameters $\lambda$ and $T$ [61]. We adjust the likelihood scores by subtracting a scaled correction term while capping its contribution at a threshold to prevent excessive influence from low-likelihood completions, ensuring a balanced adjustment across all outputs:
^$\hat{\mathcal{L}}(c_{i},(D,I,p))=\mathcal{L}(c_{i},(D,I,p))-\lambda\cdot\max(\mathcal{L}(c_{i},(\emptyset,p)),T)$ (6)
In practice, we find that ordinal scoring works better than converting raw log-likelihoods to a relevance rating. 1-5 scoring is also practically feasible to correlate with human judgment. The final CAIRe formulation uses ordinal scoring to grade images across cultures in the label set.
Model Img Wiki Top-20 F1 $\Delta$${}_{\textsc{CAIRe}}$
CAIRe
Llama-3.2-11B-Vis.-Ins. 47.4 (+6.6)
47.9 (+7.1)
[…]
52.1 (+11.1)
[…]
54.8 (+13.8)
[…]
29.0 (+8.7)
CAIRe (Log-probabilities)
Llama-3.2-11B-Vis-Ins. 61.8 (+21.0)
Qwen2.5-7B-Ins. 52.5 (+11.5)
Qwen2.5-VL-7B-Ins. 43.6 (+2.6)
Pangea-7B-hf 56.0 (+35.7)
Table 8: F1-scores on the specific set. $\Delta_{\textsc{CAIRe}}$ represents the improvement each CAIRe implementation achieves over the strongest baseline.
Appendix C Experimental Details
C.1 Culturally Similar Matches
Refer to caption Gold: Namahage CAIRe: Oni
Refer to caption Gold: Thangka Painting CAIRe: Tibetan Art
Refer to caption Gold: Day of the Dead CAIRe: Mictēcacihuātl
Figure 7: Specific set examples
There are instances where the exact entity is not matched. A few of these examples are presented in Figure 7.
Case 1: In this case, the retrieved entity represents a specific aspect of a broader cultural concept. For example, as illustrated in the first case, the retrieved entity is an Oni, a type of Japanese troll, whereas the gold-standard label, Namahage, refers to a broader category of Japanese demon-like creatures. Despite this distinction, the retrieved entity remains culturally relevant and semantically aligned with the gold label.
Case 2: The second case exemplifies instances where the retrieved entity directly represents the associated culture. The image depicts a Thangka painting, a form of Tibetan art. Our retrieval system identifies the broader concept of Tibetan art, which remains an accurate cultural attribution given the overarching goal of cultural relevance assessment.
Case 3: The final case mirrors the first, where the retrieved entity corresponds directly to the deity depicted in the image. This deity is highly relevant and closely associated with the gold-standard label, which represents the broader cultural event, the Day of the Dead.
In such cases, while the retrieval is technically classified as incorrect, the overall cultural scoring remains reliable. This is because the retrieved entities are often culturally analogous to the target entities, and their corresponding Wikipedia pages contain thematically similar content. Consequently, the models can still effectively extract the relevant cultural information.
C.2 GlobalRG Concept List
Table 9 presents the 20 human universals selected by [3].
Breakfast Clothing Dance Drinks
Dessert Dinner Farming Festival
Eating Habits Funeral Greetings Head Coverings
Instrument Lunch Marriage Music
Religion Ritual Sports Transport
Table 9: Cultural Categories Table
Model Image Wikipedia F1 Score
CAIRe (1-5 Scores)
Llama-3.2-11B-Vision-Instruct 48.8
Qwen2.5-7B-Instruct 77.7
Qwen2.5-VL-7B-Instruct 77.7
Pangea-7B-hf 46.6
CAIRe (Log-probabilities)
Llama-3.2-11B-Vision 65.7
Qwen2.5-7B 58.6
Qwen2.5-VL-7B-Instruct 45.1
Pangea-7B-hf 62.6
Table 10: F1-scores with Gold context
Model Image Ratio
CAIRe (1-5 Scores)
Llama-3.2-11B-Vision-Instruct 97.1
Qwen2.5-7B-Instruct 88.7
Qwen2.5-VL-7B-Instruct 84.3
Pangea-7B-hf 92.1
CAIRe (Log-probabilities)
Llama-3.2-11B-Vision 94.1
Qwen2.5-7B 89.6
Qwen2.5-VL-7B-Instruct 96.7
Pangea-7B-hf 89.5
Table 11: Ratio between CAIRe and Gold context
C.3 Cultural Relevance Scoring Criteria
Table 12 presents descriptions for each score in our 1-5 scoring formulation.
Score Relevance Level Description
1 Not Relevant The content does not connect with or reflect the target culture at all.
2 Minimally Relevant The content shows slight or superficial connections to the culture but lacks depth. May include vague references or isolated cultural elements that feel out of place or underdeveloped.
3 Somewhat Relevant The content contains identifiable cultural references, but they may feel generic, inconsistent, or limited in scope. The connection to the culture is present but could be stronger or more meaningful.
4 Relevant The content reflects a reasonable understanding of the culture, including accurate and appropriate references. It integrates cultural aspects well, though there may still be areas where more depth could be added.
5 Highly Relevant The content is deeply connected to the target culture, showing an immersive, accurate, and respectful understanding. Cultural references feel natural, meaningful, and central to the content.
Table 12: Cultural Relevance Scoring Criteria
C.4 Example Prompt for 1-5 Scoring
A prompt template used in the 1-5 Scoring method for VLMs is provided in Figure 8.
Prompt Template: We want to assess how relevant an image is to a given culture. We have identified this concept to be closely associated with the image: {entity}. Here is some detailed information about this concept from Wikipedia: {wiki}. Using the above context, assign a score from 1 to 5 based on how culturally relevant the image is to {target}: Think step by step, specifically considering cultural symbols, styles, traditions, or any features that align with the culture of {target}. The final score should be a number between 1 to 5, where the meaning of each score is defined as follows: 1 – Not Relevant: The content does not connect with or reflect the target culture at all. 2 – Minimally Relevant: The content shows slight or superficial connections to the culture but lacks depth. May include vague references or isolated cultural elements that feel out of place or underdeveloped. 3 – Somewhat Relevant: The content contains identifiable cultural references, but they may feel generic, inconsistent, or limited in scope. The connection to the culture is present but could be stronger or more meaningful. 4 – Relevant: The content reflects a reasonable understanding of the culture, including accurate and appropriate references. It integrates cultural aspects well, though there may still be areas where more depth could be added. 5 – Highly Relevant: The content is deeply connected to the target culture, showing an immersive, accurate, and respectful understanding. Cultural references feel natural, meaningful, and central to the content. The output should be a single number ONLY.
Example Conversation Format: System: You are an expert in evaluating the cultural relevance of images. User: [Image] [Text: Prompt Template] Model: Final Score: [1-5]
Figure 8: Prompt Template for CAIRe methods
C.5 Upper Bound Analysis
We obtain the results in Table 10 by replacing the retrieved Wikipedia pages with the ground-truth pages corresponding to the entities in our specific set. By comparing the F1-scores of our retrieved pages to those of the gold-standard pages, we quantify the retrieval-induced error in our pipeline. Specifically, we compute the ratio of the F1-scores from our VEL system to those obtained using the gold-standard pages. As shown in Table 11, our method consistently achieves scores comparable to the gold standard, highlighting the effectiveness of our retrieval system in identifying culturally relevant entities.
C.6 Specific Set Label Set
Table 13 shows the full set of candidate culture labels for the images in the Specific Set:
Category Labels
Geographical Entities Countries of the world
Countries of Africa
States of Mexico
States of the US
States of India
Major cities of India
Cities of Indonesia
Cities of Nepal
Ethnic & Cultural Groups Ethnicities of Africa
Ethnicities of India
Ethnicities of Australia
Ethnicities of the US
Ethnicities of Indonesia
Festivals & Traditions Festivals of India
Philosophy & Religion Japanese philosophy
World religions
Historical Civilizations Bronze Age civilizations
Table 13: Cultural labels for the Specific Set
C.7 Alternate Thresholds
Method Threshold=2 Threshold=3 Threshold=4 Threshold=5
F1 P R F1 P R F1 P R F1 P R
CAIRE
Qwen-LM 0.33 0.17 0.96 0.40 0.33 0.88 0.69 0.74 0.76 0.59 0.73 0.55
Qwen 0.52 0.47 0.84 0.63 0.61 0.79 0.66 0.74 0.65 0.20 0.25 0.18
Pangea 0.38 0.29 0.91 0.38 0.29 0.91 0.43 0.34 0.90 0.32 0.40 0.32
LLaMA 0.34 0.31 0.82 0.48 0.47 0.76 0.47 0.51 0.60 0.00 0.00 0.00
Baseline
Qwen 0.33 0.26 0.89 0.38 0.32 0.82 0.41 0.38 0.67 0.17 0.20 0.17
Pangea 0.19 0.12 0.96 0.19 0.12 0.96 0.20 0.13 0.96 0.07 0.08 0.09
LLaMA 0.12 0.07 1.00 0.29 0.20 0.95 0.41 0.39 0.75 0.00 0.00 0.00
Table 14: Performance across thresholds for CAIRE and Baseline methods using various VLMs.